OpenStudy (anonymous):

Your kind aunt tells you that you can choose one of the following: Option 1: During a period of 5 days, she will give you 3 cents the first day, give you double that the second day, give you double the second on the third day and so on. Option 2: She will give you $10. Which option would give you more money, 1 or 2? 1 2

jimthompson5910 (jim_thompson5910):

Option 1: Day 1: 3 cents Day 2: 2*3 = 6 cents Day 3: 2*6 = 12 cents Day 4: 2*12 = 24 cents Day 5: 2*24 = 48 cents So over the course of 5 days, option 1 will give you 3+6+12+24+48 = 93 cents This is close to $1, but nowhere near $10. It's far less. So option 2 is a much better option.